OpenStudy (anonymous):

The voltage, v, (volts) in an alternating current at any time, t, (seconds) is given by: V = 100 sin (200πt + π/4) Using differential calculus, calculate: 1. The time to reach the greatest voltage rate of change 2. The value of the greatest voltage rate of change 3. Using integral Calclulus, calculate the RMS value of the voltage

OpenStudy (cwrw238):

differentiate dV/dt and equate to zero

OpenStudy (kainui):

So first off, what is the voltage rate of change? That's the derivative of the voltage with respect to time. Now where is that at it's maximum? Well when it's at its maximum, it will have a slope of 0, so take the derivative again and set it equal to zero, and you can solve for time t. Since you're looking for the time to reach it the first time, you choose the value of t which is the smallest and non-negative, unless you have a time machine. So you differentiate twice, not once like cwrw is saying.

OpenStudy (cwrw238):

lol time machine !! yes u r right kalnui - however i was only indicating how to start the problem

OpenStudy (kainui):

Err, sorry I might have misread what you wrote. Anyways, I'm gone so good luck guys!

OpenStudy (anonymous):

So far I have calculated the amplitude =100V, Frequency = 100Hz and Phase Angle = 45degrees (don't know how to get the little zero for degrees on keyboard.) The voltage when t=0 = v=100Sin(0+0.785) = 100Sin(45Degrees) = 71V The Voltage when t=5ms = 100Sin(225degress) = -71V The time when voltage first reaches 50V.

OpenStudy (anonymous):

When V=50V then 50V= 100Sin(200*pi*t+0.785) 50/100=Sin(200*pi*t+0.785) (200*pi*t+0.785)=arcsin 50/100 =30degrees or 0.524rad 200*pi*t= 0.524+0.785 = 1.309 So when V=50v Time t = 1.309/200*pi = 2.08ms

OpenStudy (anonymous):

v=100Sin(200*pi*t+pi/4) Using differential calculus, calculate: The time for the the voltage to first reach a maximum: Maximum volatge = 100V Hence 100=100Sin(200*pi*t+0.785) 1=Sin(200*pi*t+0.785) arcsin1=200*pi*t+0.785 =90 degress or 1.5708rad 200*pi*t=1.5708 - 0.785 = 0.7858 t=0.7858/200*pi = 1.25ms

OpenStudy (anonymous):

The time to reach the greatest voltage rate of change dv/dt is a maximum when Cos(200πt+π/4) is either 1 or -1 which occurs when 200πt+π/4=0 or π radians t is either =(-1)/800 secs or 1/800 secs t=-0.00125 secs or t=0.00125 secs

OpenStudy (john5454):

Is this correct please as I have the same question